A person invested $2,000, part at 7% and part at 8% simple interest. If the total yearly interest from these investments was $154, how much was vested at each rate?


How many gallons of a 70% alcohol solution must be mixed with 30 gallons of a 15% alcohol solution to obtain a solution that is 60% alcohol?

Answer by ichudov(499)

You should ask two questions separately, but here goes.

Denote the amount invested at 7% as x dollars. The remaining amount to be invested is, obviously, 2000-x. Amount of interest at 7% is x*0.07. Amount of interest invested at 8% is (2000-x)*0.08. So, we have x*0.07+(2000-x)*0.08 = 154.

x*0.07+2000*0.08 - x*0.08 = 154

2000*0.08-154 + x*0.07 - x*0.08 = 0

160-154-0.01x = 0

6 = 0.01x

x = 600, 2000-x = 1400

Answer: He invested $600 at 7% and $1,400 at 8%.

Regarding your unhealthy interest in alcohol solutions :), I have a word problem solver for exactly this problem. Go to word problems from my home page.